[Brahms-l] Shift report 20050622 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jun 22 2005 - 15:52:34 EDT
Shifter leader was eb
Shifters were rd


1) 2.5 hours beam time - data runs 15740 thru 15743.
2) errors in displayed magnet settings, beam conditions for runs 15742, 15743 noted due to RHIC "controls failure".
3) no detector problems.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jun 22 15:53:08 2005

This archive was generated by hypermail 2.1.8 : Wed Jun 22 2005 - 15:53:21 EDT